### v4

@@ 6 Conclusion and future work
This paper presents an $O(n\log n)$-time deterministic algorithm for the planar two-center problem, which improves the previous $O(n\log^{2}n)$-time algorithm by Wang [29] and matches the $\Omega(n\log n)$ lower bound for the problem given by Eppstein [15]. Our algorithm is conceptually simple and elegant, while its analysis is technical and requires new insights to the problem. We believe that some of our techniques can find applications in other variants of planar two-center (or more generally, other geometric coverage problems).
Next, we pose some interesting open questions for future study. First, given that planar two-center can now be solved optimally, it is natural to ask for the largest $k$ such that planar $k$-center can be solved in $O(n\log n)$ time (or near-linear time). This question is completely open: even for $k=3$ with the assumption that all input points are in convex positions, the best known algorithm runs in near-quadratic time [5]. Furthermore, one can also study the two-center problem in higher dimensions and ask whether the problem can be solved in near-linear time in any fixed dimension (which is true for the one-center problem). Finally, we notice that, while near-linear time algorithms for planar two-center have been known for more than two decades since the work of Sharir [27], the best known algorithms for discrete planar two-center run in $O(n^{4/3}\cdot\text{poly}(\log n))$ time [3, 30], as mentioned in the introduction. Therefore, it is interesting to ask whether the techniques in this paper can also be applied to design near-linear time algorithms for the discrete problem.

### v5

@@ 6 Conclusion and future work
This paper presents an $O(n\log n)$-time deterministic algorithm for the planar two-center problem, which improves the previous $O(n\log^{2}n)$-time algorithm by Wang [wang2022planar] and matches the $\Omega(n\log n)$ lower bound for the problem given by Eppstein [eppstein1997faster]. Our algorithm has a straightforward design while its analysis is technically involved and requires new approaches to the problem. We believe that some of our techniques can find applications in other variants of the planar two-center problem (or more generally, other geometric coverage problems).
Next, we pose some interesting open questions for future study. First, given that the planar two-center problem can now be solved optimally, it is natural to ask for the largest $k$ such that planar $k$-center can be solved in $O(n\log n)$ time (or near-linear time). This question is completely open: even for $k=3$ with the assumption that all input points are in convex positions, the best-known algorithm runs in near-quadratic time [bian2018efficient]. Furthermore, one can also study the two-center problem in higher dimensions and ask whether the problem can be solved in near-linear time in any fixed dimension (which is true for the one-center problem). Finally, we notice that, while near-linear time algorithms for the planar two-center problem have been known for more than two decades since the work of Sharir [Sharir-twocenter], the best known algorithms for the discrete planar two-center problem run in $O(n^{4/3}\cdot\text{poly}(\log n))$ time [AgarwalPK1998TD2P, ref:WangUn23], as mentioned in the introduction. Therefore, it is interesting to ask whether the techniques in this paper can also be applied to design near-linear time algorithms for the discrete problem.
[…]